# Towards a Universal Image Degradation Model via Content-Degradation Disentanglement

### v1

@@ Abstract
Image degradation synthesis is highly desirable in a wide variety of applications ranging from image restoration to simulating artistic effects. Existing models are designed to generate one specific or a narrow set of degradations, which often require user-provided degradation parameters. As a result, they lack the generalizability to synthesize degradations beyond their initial design or adapt to other applications. Here we propose the first universal degradation model that can synthesize a broad spectrum of complex and realistic degradations containing both homogeneous (global) and inhomogeneous (spatially varying) components. Our model automatically extracts and disentangles homogeneous and inhomogeneous degradation features, which are later used for degradation synthesis without user intervention. A disentangle-by-compression method is proposed to separate degradation information from images. Two novel modules for extracting and incorporating inhomogeneous degradations are created to model inhomogeneous components in complex degradations. We demonstrate the model’s accuracy and adaptability in film-grain simulation and blind image restoration tasks. The demo video, code, and dataset of this project will be released upon publication at https://github.com/yangwenbo99/content-degradation-disentanglement.
@@ 1 Introduction
Therefore, it is desirable to design a universal degradation model that can generalize to complex (combinations of) degradations, both homogeneous and inhomogeneous. However, achieving this is challenging for two main reasons. First, complex degradations are harder to separate from degraded images than simpler ones. Simple degradation models often assume linear or affine, content-independent degradation processes. However, real-world complex degradations are typically nonlinear and content-dependent, complicating the extraction of pure degradation information free of image content. Second, extracting and synthesizing inhomogeneous degradations presents additional challenges, as these spatially varying effects are more difficult to model than homogeneous ones.
@@ 2 Related Works
Table 1 summarizes representative degradation models. Most existing models focus on a single distortion type, while some handle basic fixed combinations of distortions. For instance, the super-resolution model [44] considers common degradations in the imaging pipeline: blur, downsampling, noise and compression artifacts. However, each degradation component is treated independently. Some components are handcrafted, requiring storng domain knowledge specific to the downstream task (super-resolution in this case). Similarly, the degradation model in another super-resolution method [35] simulates the imaging pipeline twice using a fixed sequence of degradations. These degradation models cannot be easily transferred to other degradation types.
@@
Another related area is Style Transfer (ST), which aims to impose the artistic style of a source image to the content of a target image. However, most ST methods require training a separate model for each specific artistic style. A few SOTA methods, such as StyTr2 [41], handle multiple styles by extracting style information from a separate input image [22, 9]. Nevertheless, image degradations exhibit distinct characteristics from artistic styles, necessitating specialized architecture for effective modeling. As a result, SOTA ST models (e.g., StyTr2 [41]) exhibit significant limitations (Fig. 3): they create undesirable changes to image content (e.g., skin tone), are poor at reproducing realistic noise patterns, are incapable of segregating homogeneous and inhomogeneous degradations, and fail at handling inhomogeneous degradations due to architectural limitations.
@@ 3.2 Homo-/inhomogeneous degradation synthesis
We identify providing a smooth channel of incorporating local degradation as a practical requirement for disentangling local degradation information from global. Performing convolution using a spatially-varying kernel is an ideal way of providing such channel. However, its space complexity is prohibitively expensive. The main cost is attributed to the inter-channel dependencies, while spatial dependencies are mainly handled by depthwise components. To reduce the computation cost, we propose an inhomogeneous degradation-aware (IDA) layer:
@@ 3.3 Degradation disentanglement
where $p$ is the probability density. The density $p\left(e_{g}^{(i)}\right)$ and $p\left(e_{l}^{(i,j)}\right)$ are estimated using [7] and [6], respectively. Due to the large sample size used in the estimation of $p$, its estimation is accurate enough (while $p\left(\mathbf{e}_{g}\right)$ cannot be reliably estimated, due to its high dimensionality).
@@
The following mild assumptions are made: (A-1) With information about the degradation process and the clean image, the degraded image can be exactly reconstructed, albeit the random state difference. (A-2) The distribution of distortion processes is independent of the clean images’. (A-3) The distortion can be inferred from the degraded image. (A-4) When an appropriate perceptual similarity loss $\mathcal{L}_{\text{sim }}=d\left(\mathbf{y},\hat{\mathbf{y}}\right)$ is used, the reconstructed distorted image $\hat{\mathbf{y}}$ will be be similar enough to $\mathbf{y}$. (See supplementary materialsfor detailed justifications for each assumption.) With these assumptions, we are able to rigorously prove (in Supp. Sec. 6.2) that
@@
We further assume that (A-5) there exists two kinds of degradations: the homogeneous one $\mathbf{d}_{g}$ is the same for all geometry location, and the inhomogeneous one $\mathbf{d}_{l}^{(i)}$ is different for different geometry location $(i)$, meaning they are jointly independent. Then, we show in the supplementary materials that, with a flexible enough network, the optimal solution for $\lambda_{g}\mathcal{L}_{\text{rate\_g}}+\lambda_{l}\mathcal{L}_{\text{rate\_l}}$ guarantees that $\mathbf{e}_{l}$ only contains spatially dependent information, and $\mathbf{e}_{g}$ only contains homogeneous distortion information.
@@ 4.1 Degradation reproduction and transfer
To test our model’s generalizability to image semantics, we curated a dataset of 300K Wikimedia Quality Images (WQIs), each of which is inspected by Wikipedia or its sister projects’ editors for high technical and aesthetic quality [19]. From this, we generated 40K training pairs using about 13K images in the dataset. Each pair consists of a pristine image $\mathbf{x}$ and a distorted image $\mathbf{y}$, generated by distorting $\mathbf{x}$ with distortions randomly selected from a pool of 16 diverse degradation combinations designed according to typical image processing pipelines. Details of the dataset and training settings are in Supp. Sec. 4. Figure 7 showcases our model’s ability to reproduce and transfer various degradations (e.g., Gaussian noise, blur, motion blur, JPEG artifacts) and their combinations across different images. Unlike [12], our model generates diverse noise patterns, demonstrating both effective representation learning and output diversity.
@@
Authentic image degradations often include inhomogeneous elements that cannot be easily modeled by simple statistical models. We employed a real-world rainy image dataset [33]. As demonstrated in Figure 8, our model successfully reproduces and transfers these complex, authentic distortions. Extra tests on the GoPro [27] dataset is in Supp. Sec. 3.2.
@@ 4.2 Degradation latent disentanglement
By analyzing variances, we identified five active dimensions in $\mathbf{e}_{g}$ for the model trained on the distorted WQI dataset. The model learns to disentangle degradation components and enables human to find interpretable meaning of each active latent dimension. To demonstrate this, we modified each dimension of the degradation embedding for a pristine test image. Figure 9 shows visual results for one of the active dimensions, with additional examples available in the Supp. Sec. 3.3. Most degradation embedding dimensions control a group of degradations, aligning with our understanding of natural biases and masking effects [37] in the dataset.
4.3 Film grain encoding and transfer
@@ 4.5 Ablation Studies
Figure 12: Examples of degradation transfer from mixed sources. Without the IDA layer or IDE network (IDEN), the model is unable to properly reproduce inhomogeneous degradation. Inputs to the models ($\mathbf{x},\mathbf{y}_{1},\mathbf{y}_{2}$) are shown in Fig. 2.
@@
Since all three cases use distortions synthesized from scripts, we are able to generate ground-truth distorted images to evaluate our performance. The distorted FFHQ dataset constructed in the previous subsection is used to train the models analyzed in the ablation study.

### v2

@@ Abstract
Image degradation synthesis is highly desirable in a wide variety of applications ranging from image restoration to simulating artistic effects. Existing models are designed to generate one specific or a narrow set of degradations, which often require user-provided degradation parameters. As a result, they lack the generalizability to synthesize degradations beyond their initial design or adapt to other applications. Here we propose the first universal degradation model that can synthesize a broad spectrum of complex and realistic degradations containing both homogeneous (global) and inhomogeneous (spatially varying) components. Our model automatically extracts and disentangles homogeneous and inhomogeneous degradation features, which are later used for degradation synthesis without user intervention. A disentangle-by-compression method is proposed to separate degradation information from images. Two novel modules for extracting and incorporating inhomogeneous degradations are created to model inhomogeneous components in complex degradations. We demonstrate the model’s accuracy and adaptability in film-grain simulation and blind image restoration tasks. The demo video, code, and dataset of this project will be released at Yang’s GitHub account .
@@ 1 Introduction
Therefore, it is desirable to design a universal degradation model that can generalize to complex (combinations of) degradations, both homogeneous and inhomogeneous. However, achieving this is challenging for two main reasons. First, complex degradations are harder to separate from degraded images than simpler ones. Simple degradation models often assume linear (or affine) and content-independent degradation processes. However, real-world complex degradations are typically nonlinear and content-dependent, complicating the extraction of pure degradation information free of image content. Second, extracting and synthesizing inhomogeneous degradations presents additional challenges, as these spatially varying effects are more difficult to model than homogeneous ones.
@@ 2 Related Works
Table 1 summarizes representative degradation models. Most existing models focus on a single distortion type, while some handle basic fixed combinations of distortions. For instance, some super-resolution models [44, 36] considers common degradations in the imaging pipeline such as blur, downsampling, noise and compression artifacts. However, each degradation component is modeled independently. Some components are handcrafted, requiring strong domain knowledge specific to the downstream task (super-resolution in this case). These degradation models cannot be easily transferred to other degradation types.
@@
Another related area is Style Transfer (ST), which aims to impose the artistic style of a source image to the content of a target image. However, most ST methods require training a separate model for each specific artistic style. A few SOTA methods handle multiple styles by extracting style information from a separate input image [23, 9, 16]. Nevertheless, image degradations exhibit distinct characteristics from artistic styles, necessitating specialized architecture for effective modeling. As a result, SOTA ST models (e.g., StyTr2 [16]) exhibit significant limitations (Fig. 3): they create undesirable changes to image content (e.g., skin tone), are poor at reproducing realistic noise patterns, are incapable of segregating homogeneous and inhomogeneous degradations, and fail at handling inhomogeneous degradations due to architectural limitations.
@@ 3.2 Homo-/inhomogeneous degradation synthesis
We identify providing a smooth channel of incorporating local degradation as a practical requirement for disentangling local degradation information from global. Although performing convolution using a spatially-varying kernel is an ideal way of providing such channel, its space complexity is prohibitively expensive. The main cost is attributed to the inter-channel dependencies, while spatial dependencies are mainly handled by depthwise components. To reduce the computation cost, we propose an inhomogeneous degradation-aware (IDA) layer:
@@ 3.3 Degradation disentanglement
where $p$ denotes the probability density function. The density $p\left(e_{g}^{(i)}\right)$ and $p\left(e_{l}^{(i,j)}\right)$ are estimated using [7] and [6], respectively. Due to the large sample size used in the estimation of $p$, its estimation is accurate enough (while $p\left(\mathbf{e}_{g}\right)$ cannot be reliably estimated, due to its high dimensionality).
@@
The following mild assumptions are made: (A-1) With information about the degradation process and the clean image, the degraded image can be exactly reconstructed, albeit the random state difference. (A-2) The distribution of distortion processes is independent of the clean images’. (A-3) The distortion can be inferred from the degraded image. (A-4) When an appropriate perceptual similarity loss $\mathcal{L}_{\text{sim }}=d\left(\mathbf{y},\hat{\mathbf{y}}\right)$ is used, the reconstructed distorted image $\hat{\mathbf{y}}$ will be be similar enough to $\mathbf{y}$. (See Supp. Sec. 2.2 for detailed justifications.) With these assumptions, we are able to rigorously prove (in Supp. Sec. 6.2) that
@@
We further observe that there exists two kinds of degradations: the homogeneous one $\mathbf{d}_{g}$ is the same for all geometry location, and the inhomogeneous one $\mathbf{d}_{l}^{(i)}$ is different for different geometry location $(i)$. Based on our observations, we assume (A-5) they are jointly independent. Then, we show in the supplementary materials that, with a flexible enough network (and $\lambda_{l}$ and $\lambda_{g}$ satisfying mild conditions), the optimal solution for $\lambda_{g}\mathcal{L}_{\text{rate\_g}}+\lambda_{l}\mathcal{L}_{\text{rate\_l}}$ guarantees that $\mathbf{e}_{l}$ only contains spatially dependent information, and $\mathbf{e}_{g}$ only contains homogeneous distortion information.
@@ 4.1 Degradation reproduction and transfer
To test our model’s generalizability to image semantics, we curated a dataset of 300K Wikimedia Quality Images (WQIs), each of which is inspected by Wikipedia or its sister projects’ editors for high technical and aesthetic quality [20]. From this, we generated 40K training pairs using about 13K images in the dataset. Each pair consists of a pristine image $\mathbf{x}$ and a distorted image $\mathbf{y}$, generated by distorting $\mathbf{x}$ with distortions randomly selected from a pool of degradation combinations designed according to typical image processing pipelines. Details of the dataset and training settings are in Supp. Sec. 4. Figure 7 showcases our model’s ability to reproduce and transfer various degradations (e.g., Gaussian noise, blur, motion blur, JPEG artifacts) and their combinations across different images. Unlike [12], our model generates diverse noise patterns, demonstrating both effective representation learning and output diversity.
@@
By analyzing variances, we identified five active dimensions in $\mathbf{e}_{g}$ for the model trained on the distorted WQI dataset. The model learns to disentangle degradation components and enables human to find interpretable meaning of each active latent dimension.
Authentic image degradations often include inhomogeneous elements that cannot be easily modeled by simple statistical models; for example, real-world raining effects [34]. As shown in Figure 8, our model successfully reproduces and transfers these complex, authentic distortions. Extra tests on the GoPro dataset [28] is in Supp. Sec. 3.2.
@@ 4.2 Degradation latent disentanglement
For demonstration, we modified each active dimension of the degradation embedding for a pristine test image. Figure 9 shows visual results for one of the active dimensions, with additional examples available in the Supp. Sec. 3.3. Most degradation embedding dimensions control a group of degradations, aligning with our understanding of natural biases and masking effects [38] in the dataset.
4.3 Film grain encoding and transfer
@@ 4.5 Ablation Studies
Figure 12: Examples of degradation transfer from mixed sources. Without the IDA layer or IDEN, the model is unable to properly reproduce inhomogeneous degradation. Inputs to the models ($\mathbf{x},\mathbf{y}_{1},\mathbf{y}_{2}$) are shown in Fig. 2.
@@
Since all three cases use distortions synthesized from Python scripts, we are able to generate ground-truth distorted images to evaluate our performance. The distorted FFHQ dataset constructed in the previous subsection is used to train the models analyzed in the ablation study.
@@ 5 Conclusion
Acknowledgements
This work is supported in part by National Sciences and Engineering Research Council of Canada under the Canada Research Chair and Discovery Grant programs.